OpenStudy (anonymous):

a. y = 3(x + 8) solve for x

OpenStudy (anonymous):

First, distribute the 3. y=3x+24 Then move the 24 to the other side by subtracting. y-24=3x Now divide by 3. 1/3y-8=x

OpenStudy (anonymous):

wait so would that be the answer..?

OpenStudy (anonymous):

x=y/3-8 Do you understand the steps?

OpenStudy (anonymous):

not reall, i dont understand what it means b "solve for x"

OpenStudy (anonymous):

Solving for X simply means that you want to have X by itself on on side of the equation. So solve for x: 2-x=y

OpenStudy (anonymous):

y plus two?

OpenStudy (anonymous):

close! 2-x=y (we need to get that x by itself) add the X to both sides 2=y+x Now subtract the y from both sides 2-y=x So the answer is 2-y or x=2-y

OpenStudy (anonymous):

ohhhh!

OpenStudy (anonymous):

Want to try one more time?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Solve for X: 5x=y

OpenStudy (anonymous):

y/5 ?

OpenStudy (anonymous):

Correct!!! Usually teachers prefer that you answer x=y/5 though.. But its still correct!

OpenStudy (anonymous):

yay! thanks sooooo much!

OpenStudy (anonymous):

No problem!:)